# Physics-Informed Detection of Friction Anomalies in Satellite Reaction Wheels

### v1

@@ IV-C Maximum Likelihood Assignment
Once the total dry friction coefficient $\hat{f}^{\mathrm{d}}_{i}$ has been estimated for each interval, we are interested in decomposing this friction among the different FSSs, that is, to obtain the value of the various components of (4b) at every step, rather than only the sum of them. To achieve this, we use the knowledge of the transition probability of the subsystems defined in (5) together with the steps $k_{i}^{\mathrm{cp}}$ at which the friction changes took place to determine the most likely sequence of configurations that the subsystems went through.
Once this sequence has been estimated, we can determine the values of each subsystem’s friction coefficients $\hat{f}_{k}^{s}$ and the base dry friction $\hat{f}^{\bar{\mathrm{d}}}$. Since the dry and viscous coefficients are a function of the anomalies present in the RWA, as considered in Equation (9), we will use their values for the diagnosis.
Formally, the algorithm takes the friction values, rejection costs, and location of the changepoints as input. It estimates the base dry friction, and the friction values of each subsystem at each configuration (that is, all the elements in (4b)).
@@ IV-C1 Maximum likelihood jumping
Extending the notation used in Section III-B, we define the state of the RWA at every interval $i$ with a tuple of two elements: the configuration $\mathbf{q}_{i}$ of each subsystem and the number of steps $\boldsymbol{\tau}_{i}$ since each subsystem changed configuration for the last time, both given as vectors with as many elements as subsystems present in the RWA, so $x_{i}=(\mathbf{q}_{i},\boldsymbol{\tau}_{i})$, with $\mathbf{q}_{i}=\begin{bmatrix}q_{i}^{1}&\cdots&q_{i}^{N_{\mathrm{s}}}\end{bmatrix}^{T}$ and $\boldsymbol{\tau}_{i}=\begin{bmatrix}\tau_{i}^{1}&\cdots&\tau_{i}^{N_{\mathrm{s}}}\end{bmatrix}^{T}$.
@@
With this notation, $q_{i}^{s}$ represents the configuration of the FSS with index $s$ at interval $i$. Every changepoint $i$ is characterized by two variables: the number of steps since the previous changepoint ($\Delta\tau_{i}$), and the sign of the change of friction ($\Delta f_{i}$), which can be either 1-1 for a decrease or $1$ for an increase of the friction. At a changepoint, one of the subsystems changes its configuration by $\Delta f_{i}$ and its number of steps since the last changepoint is reset to 0. The other subsystems have unaffected configurations, but their $\tau$ increases by $\Delta\tau_{i}$. We choose which subsystem has jumped at each stage with the input $u_{i}$. The set of possible inputs is $\mathcal{U}\in\llbracket 1,N_{\textrm{s}}\rrbracket$, and the transition function is, according to the description given before,
$h_{i}(\mathbf{x}_{i},u_{i})$ $=\left(\mathbf{q}_{i}+\Delta f_{i}\mathbf{e}_{u_{i}},(\boldsymbol{\tau}_{i}+\Delta\tau_{i})(\mathbb{1}-\mathbf{e}_{u_{i}})\right),$ (19)
@@
where $\mathbb{1}$ is a vector with all entries equal to $1$ and $\mathbf{e}_{u}$ is a vector with all entries 0 except the $u$-th one, which takes a value of $1$. A visualization of the possible transitions of a system is given in Figure 6.
@@
Regarding the transition costs of the problem, they are given by the log-likelihood of the event:
$g_{i}(x_{i},u_{i})$ $=\sum_{s\not=u_{i}}\log P_{s}^{\neg\mathrm{t}}(q_{s},\tau_{s},\tau_{s}+\Delta\tau)$
$+$ $\log P_{u_{i}}^{\neg\mathrm{t}}(q_{u_{i}},\tau_{u_{i}},\tau_{u_{i}}+\Delta\tau_{u_{i}}-1)$ (20)
$+$ $\log P_{u_{i}}^{\mathrm{t}}(q_{u_{i}},\tau_{u_{i}}+\Delta\tau)+\log P_{u_{i}}^{\mathrm{q}}(q_{u_{i}}+\Delta f_{i}|q_{u_{i}}).$
In this computation, we define $P_{s}^{\neg t}(q,k_{1},k_{2})$ as the probability of the switching system $s$ not transitioning away from its configuration $q$ between steps $k_{1}$ and $k_{2}$. It can be calculated as
$P_{s}^{\neg\mathrm{t}}(q,k_{1},k_{2})=\sum_{k=k_{1}}^{k_{2}}\log\left[1-P_{s}^{\mathrm{t}}(q,k)\right].$ (21)
@@
Recalling Remark 1, we know $P_{s}^{\mathrm{t}}(q,\tau)$ and $P_{s}^{\mathrm{q}}(q^{\prime}|q)$, so we can compute all the log-likelihood scores, either as the dynamic programming problem is being solved, or in advance, and then use look-up tables during the execution of the algorithm. The objective, then, is to obtain a sequence of inputs $u$ maximizes log-likelihood. This is a standard dynamic programming problem that can be solved with classical algorithms [46]. Note that while the possible values of $\mathbf{q}$ are relatively limited, this is not the case for $\boldsymbol{\tau}$, and the number of possible states may increase exponentially with the number of stages.
@@
For the initial state $x_{0}$ of the dynamic programming problem, we take an optimistic approach: for $\mathbf{q}_{0}$, we initialize the algorithm considering all possible values, and keep the one that provides the maximum final score. Regarding $\boldsymbol{\tau}_{0}$, we select the value that maximizes the score in (20).
@@
The solution $u^{*}$, obtained by solving (18), provides then the index of FSS that produced each sudden change of friction observed in the data. Additionally, we denote $\mathbf{x}^{*}$ the sequence of states obtained by applying this input.
$\mathbf{q}_{i}$ $=\begin{bmatrix}0\\ 1\end{bmatrix}$ $\boldsymbol{\tau}_{i}$ $=\begin{bmatrix}50\\ 0\end{bmatrix}$ $\mathbf{q}_{i+1}=\begin{bmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}\\ 1\end{bmatrix},\boldsymbol{\tau}_{i+1}=\begin{bmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}\\ 80\end{bmatrix}$ $\mathbf{q}_{i+1}=\begin{bmatrix}0\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}\end{bmatrix},\boldsymbol{\tau}_{i+1}=\begin{bmatrix}130\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}\end{bmatrix}$ $\mathbf{q}_{i+1}=\begin{bmatrix}0\\ 1\end{bmatrix},\boldsymbol{\tau}_{i+1}=\begin{bmatrix}130\\ 80\end{bmatrix}$ $u_{i}=1$$u_{i}=2$$u_{i}=0$
@@ IV-C2 Changepoint rejection
The algorithm presented is inherently fragile: if a changepoint is missed during the first stage of the algorithm, or a false positive is obtained, the algorithm will fail to obtain the correct sequence of configurations of the FSSs. Unfortunately, designing a changepoint detection algorithm that detects all actual changepoints with no possibility of false positives is impossible due to the randomness of the noise that affects the friction torque measurements.
Since we want the algorithm to perform well even in the presence of false positives, we robustify it by including in the dynamic programming problem the option at every stage of ignoring the changepoint detected, rather than fitting it in one of the FSS. This is done by including 0 as an additional possible input, shown in Figure 6, with the following transition and cost functions:
@@
$g(x_{i},0)$ $=c_{i}^{\text{rjct}}+\sum_{s}P^{\neg\mathrm{t}}(q_{s},\tau_{s},\tau_{s}+\Delta\tau).$ (23)
@@ IV-C3 Friction Values
Once the dynamic programming algorithm is finished, we get a vector $u$ containing the index of the FSS that changed configuration at each changepoint and an array $\mathbf{q}$ with the configuration of every FSS after every changepoint. The next objective is to estimate the base dry friction $\hat{f}^{\bar{\mathrm{d}}}$ and, for each FSS, the sequence of configurations and friction values it goes through. We denote them as $\mathcal{Q}_{s}$ and $\mathcal{F}_{s}$ respectively, both defined as lists of values in the programming sense, where $s$ denotes the index of the FSS. At every interval, the total dry friction torque must be the base dry friction plus the contribution of each FSS as given by (4b).
This problem is solved in Algorithm 1. We start by assuming that, before the first changepoints, the friction produced by each FSS is equal to 0. Then, we compute the following friction values from the differences obtained at each changepoint. Finally, since the minimum value produced by an FSS is assumed to be 0, we adjust the coefficients to get the final values.
@@
The final output of this stage is a tuple of four elements: the estimated base dry friction ($\hat{f}^{\bar{d}}$, the estimated viscous friction coefficient ($\hat{f}^{v}$), and the sequence of states ($\mathcal{Q}_{s}$) and friction values ($\mathcal{F}_{s}$) each FSS goes through: $(\hat{f}^{\bar{\mathrm{d}}},\hat{f}^{\mathrm{v}},\{\mathcal{Q}_{s},\mathcal{F}_{s}\}_{s=1}^{N_{\mathrm{s}}})$.
@@ IV-D Anomaly Classification
$\hat{\theta}(h)=\text{sign}\left(w^{T}z+b\right).$ (26)
The classifier is parametrized by the weights $w$ and the bias $b$. Although not as versatile as state-of-the-art neural networks, SVMs have two important advantages due to their reduced number of parameters: they are less prone to overfitting, and doing inference is computationally less expensive. Although we use SVMs to classify all anomalies, the implementations vary depending on whether we are determining a dry friction anomaly, a viscous friction anomaly, or an anomaly related to an FSS.
@@ IV-D1 Dry and viscous friction anomalies
The first two anomalies involve an increased value in the base dry and viscous friction coefficients, respectively. The classification, therefore, consists in analyzing a single value in each case, either $z=\hat{f}^{\bar{\mathrm{d}}}$ or $z=\hat{f}^{\mathrm{v}}$.
This is accomplished by defining an SVM that takes values in $\mathbb{R}^{1}$ as inputs, making the classification a simple threshold learned from the dataset.
@@ IV-D2 Subsystem anomalies
Classifying anomalies for switching systems is more challenging, since we do not have a single value to classify with, but a sequence of them in the form of $\mathcal{F}_{s}$. As mentioned in Subsection III-C, an anomaly impacts the probability distribution of the friction produced by the FSS, so constructing a histogram from the friction values and feeding it to an SVM that does the classification is a reasonable approach.
The normalized histogram vector of the FSS with index $s$ is defined as $z\in\mathbb{R}^{n_{\text{bins}}}$, obtained from a sequence of friction values $f_{k}$ as
@@
Two hyperparameters must be defined, the number of bins ($n_{\text{bins}}$) and the limit values ($r_{\text{min}}$ and $r_{\text{max}}$). The latter can be determined from the historical data, by annotating the minimum and maximum values of friction observed. The number of bins to select is a trade-off between accuracy and complexity. A small number of bins may make distinguishing an anomaly impossible, while a large number will lead to a large histogram vector, which may complicate training and lead to overfitting. For our task, we select a value of $20$ bins, and later analyze the effect of this hyperparameter on the accuracy of the algorithm in Section V-D2.
@@
The last decision to take is which values of friction to include in the histogram. From the previous section, we recall that we obtain, for each subsystem, a collection of friction values paired with their corresponding configurations. Based on our knowledge of the anomaly, we can decide to include all friction values produced by the FSS in the histogram, or only those related to specific configurations (for example, only friction values obtained with $q=1$).
@@
The output of this final stage is a tuple with the status of each possible anomaly, $(\hat{\theta}^{\mathrm{d}},\hat{\theta}^{\mathrm{v}},\{\hat{\theta}^{s}\}_{s=1}^{N_{\mathrm{s}}})$.
V Results
We now proceed to analyze the model and the algorithm developed in the paper. As mentioned in Section I, this work aims to create an anomaly detection algorithm for the TAS use case.
There are three objectives pursued in this section. The first is to show the adequacy of the model proposed in Section III to the dataset provided by TAS, so we can confirm that it captures the behavior observed in the dataset with accuracy. The second objective is to analyze the algorithm’s effectiveness in detecting anomalies for the TAS use case. Finally, we want to study the effect of some hyperparameters of the algorithm whose tuning is not straightforward.

### v2

@@ IV-C Maximum Likelihood Assignment
Once the total dry friction coefficient $\hat{f}^{\mathrm{d}}_{i}$ has been estimated for each interval, we are interested in decomposing this friction among the FSSs, that is, to obtain the value of the various components of (5) at every step, rather than only their sum. To achieve this, we use the knowledge of the transition probability of the subsystems $P_{s}^{\mathrm{q}}(q|q^{\prime})$ and $P_{s}^{\tau}(\tau|q)$, together with the steps $k_{i}^{\mathrm{cp}}$ at which the friction changes took place to determine the most likely sequence of configurations that the subsystems went through.
Once this sequence has been estimated, we determine the values of each subsystem’s friction coefficients $\hat{f}_{k}^{s}$ and the base dry friction $\hat{f}^{\bar{\mathrm{d}}}$. Since the dry and viscous coefficients are a function of the anomalies present in the RWA, as in (9), we use their values for the diagnosis.
Formally, the algorithm takes the friction values, rejection costs, and location of the changepoints as input. It estimates the base dry friction, and the friction values of each subsystem at each configuration (that is, all the values in (5)).
@@ IV-C1 Maximum likelihood jumping
Extending the notation used in Section III-B, we define the state of the RWA at every interval $i$ with a tuple of two elements: the configuration $\bar{q}_{i}$ of each subsystem and the number of steps $\bar{\tau}_{i}$ since each subsystem changed configuration for the last time, both given as vectors with as many elements as subsystems present in the RWA, so $x_{i}=(\bar{q}_{i},\bar{\tau}_{i})$, with $\bar{q}_{i}=\begin{bmatrix}q_{i}^{1}&\cdots&q_{i}^{N_{\mathrm{s}}}\end{bmatrix}^{T}$ and $\bar{\tau}_{i}=\begin{bmatrix}\tau_{i}^{1}&\cdots&\tau_{i}^{N_{\mathrm{s}}}\end{bmatrix}^{T}$.
@@
With this notation, $q_{i}^{s}$ represents the configuration of the FSS with index $s$ at interval $i$. Every changepoint $i$ is characterized by two variables: the number of steps since the previous changepoint ($\Delta\tau_{i}$), and the sign of the friction change ($\Delta f_{i}$), which can be either 1-1 for a decrease or $1$ for an increase of the friction. At a changepoint, one of the subsystems changes its configuration by $\Delta f_{i}$ and its number of steps since the last changepoint is reset to 0. The other subsystems have unaffected configurations, but their $\tau$ increases by $\Delta\tau_{i}$. We choose which subsystem has jumped at each stage with the input $u_{i}$. The set of possible inputs is $\mathcal{U}\in[1,N_{\textrm{s}}]$, and the transition function is, according to the description given before,
$h_{i}(\bar{x}_{i},u_{i})$ $=\left(\bar{q}_{i}+\Delta f_{i}\bar{e}_{u_{i}},(\bar{\tau}_{i}+\Delta\tau_{i})(\mathbb{1}-\bar{e}_{u_{i}})\right),$ (26)
@@
where $\mathbb{1}$ is a vector with all entries equal to $1$ and $\bar{e}_{u}$ is a vector with all entries 0 except the $u$-th one, equal to $1$. A visualization of these transitions is shown in Figure 6.
@@
Regarding the transition costs of the problem, they are given by the log-likelihood of the transitions of the individual FSS:
$g_{i}(x_{i},u_{i})$ $=\sum_{s\not=u_{i}}\log P_{s}^{\tau}(\tau\geq\tau_{s}+\Delta\tau|q,\tau\geq\tau_{s})$
$+$ $\log P_{u_{i}}^{\tau}(\tau=\tau_{u_{i}}+\Delta\tau_{i}|q_{u_{i}},\tau\geq\tau_{u_{i}})$ (27)
$+$ $\log P_{u_{i}}^{\mathrm{q}}(q_{u_{i}}+\Delta f_{i}|q_{u_{i}}).$
In this computation, conditional probabilities of $\tau$ are computed as $P_{s}^{\tau}(\tau\geq\tau_{1}|q,\tau_{2})=P(\tau\geq\tau_{1}|q)/P(\tau\geq\tau_{2}|q)$ when $\tau_{1}\geq\tau_{2}$. In a similar way, $P_{s}^{\tau}(\tau=\tau_{1}|q,\tau_{2})=P(\tau=\tau_{1}|q)/P(\tau\geq\tau_{2}|q)$ under the same conditions.
@@
Recalling Remark 1, we know $P_{s}^{\tau}(q,\tau)$ and $P_{s}^{\mathrm{q}}(q^{\prime}|q)$, so we can compute all the log-likelihood scores, either as the dynamic programming problem is being solved, or in advance, and then use look-up tables during the execution of the algorithm. The objective, then, is to obtain an input sequence $u$ that maximizes the log-likelihood. This is a standard dynamic programming problem that can be solved with classical algorithms [5]. Note that while the possible values of $\bar{q}$ are relatively limited, this is not the case for $\bar{\tau}$, and the number of possible states may increase exponentially with the number of stages.
@@
For the initial state $x_{0}$ of the dynamic programming problem, we take an optimistic approach: for $\bar{q}_{0}$, we initialize the algorithm considering all possible values, and keep the one that provides the maximum final score. Regarding $\bar{\tau}_{0}$, we select the value that maximizes the score in (27).
@@
The solution $u^{*}$, obtained by solving (25), provides then the index of the FSS that produced each sudden change of friction observed in the data. Additionally, we denote with $\bar{x}^{*}$ the sequence of states obtained by applying this input.
$\mathbf{q}_{i}$ $=\begin{bmatrix}0\\ 1\end{bmatrix}$ $\bar{\tau}_{i}$ $=\begin{bmatrix}50\\ 0\end{bmatrix}$ $\bar{q}_{i+1}=\begin{bmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}\\ 1\end{bmatrix},\bar{\tau}_{i+1}=\begin{bmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}\\ 80\end{bmatrix}$ $\bar{q}_{i+1}=\begin{bmatrix}0\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}\end{bmatrix},\bar{\tau}_{i+1}=\begin{bmatrix}130\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}\end{bmatrix}$ $\bar{q}_{i+1}=\begin{bmatrix}0\\ 1\end{bmatrix},\bar{\tau}_{i+1}=\begin{bmatrix}130\\ 80\end{bmatrix}$ $u_{i}=1$$u_{i}=2$$u_{i}=0$
@@ IV-C2 Changepoint rejection
The algorithm presented is inherently fragile: if a changepoint is missed during the first stage of the algorithm, or a false positive is obtained, the algorithm will fail to obtain the correct sequence of configurations of the FSSs. Since we want the algorithm to perform well even in the presence of false positives, we robustify it by including in the dynamic programming problem the option at every stage of ignoring the changepoint detected, rather than fitting it in one of the FSSs. This is done by including 0 as an additional possible input, shown in Figure 6, with the following transition and cost functions:
@@
$g(x_{i},0)$ $=c_{i}^{\text{rjct}}+\sum_{s}P^{\tau}(\tau\geq\tau_{s}+\Delta\tau_{i}|q_{s},\tau\geq\tau_{s}).$ (29)
@@ IV-C3 Friction Values
Once the dynamic programming algorithm is finished, we get a vector $u$ containing the index of the FSS that changed configuration at each changepoint and an array $\bar{q}$ with the configuration of every FSS after every changepoint. The next objective is to estimate the base dry friction $\hat{f}^{\bar{\mathrm{d}}}$ and, for each FSS, the sequence of configurations and friction values it goes through. We denote them as $\mathcal{Q}_{s}$ and $\mathcal{F}_{s}$ respectively, both defined as lists of values in the programming sense, where $s$ is the index of the FSS. At every interval, the total dry friction torque must be the base dry friction plus the contribution of each FSS as given by (5).
This problem is solved in Algorithm 1. We first assume that, before the first changepoint, the friction produced by each FSS is equal to 0. Then, we compute the following friction values from the differences obtained at each changepoint. Finally, since the minimum value produced by an FSS is assumed to be 0, we adjust the coefficients to get the final values.
@@
The output of this stage is a tuple of four elements: the estimated base dry friction ($\hat{f}^{\bar{d}}$, the estimated viscous friction coefficient ($\hat{f}^{v}$), and the sequence of states ($\mathcal{Q}_{s}$) and friction values ($\mathcal{F}_{s}$) each FSS goes through: $(\hat{f}^{\bar{\mathrm{d}}},\hat{f}^{\mathrm{v}},\{\mathcal{Q}_{s},\mathcal{F}_{s}\}_{s=1}^{N_{\mathrm{s}}})$.
@@ IV-D Anomaly Classification
$\hat{\theta}(z)=\text{sign}\left(w^{T}z+b\right).$ (32)
The classifier is parametrized by the weights $w$ and the bias $b$. Although not as versatile as state-of-the-art neural networks, SVMs have two important advantages due to their reduced number of parameters: they are less prone to overfitting, and doing inference is computationally less expensive. Although we use SVMs to classify all anomalies, the implementations vary depending on whether we are determining a dry/viscous friction anomaly, or an anomaly related to an FSS.
@@ IV-D1 Dry and viscous friction anomalies
The first two anomalies involve an increased value in the base dry and viscous friction coefficients, respectively. The classification, therefore, consists in analyzing a single value in each case, either $z=\hat{f}^{\bar{\mathrm{d}}}$ or $z=\hat{f}^{\mathrm{v}}$. This is accomplished by defining an SVM that takes values in $\mathbb{R}^{1}$ as inputs, making the classification a simple threshold learned from the dataset.
@@ IV-D2 Subsystem anomalies
Classifying anomalies for switching systems is more challenging, since we do not have a single value to classify with, but a sequence of them in the form of $\mathcal{F}_{s}$. As mentioned in Subsection III-C, an anomaly impacts the probability distribution of the friction produced by the FSS, so creating a histogram of the values obtained and using an SVM that does the classification is a reasonable approach.
The normalized histogram vector of the FSS with index $s$ is defined as $z\in\mathbb{R}^{n_{\text{bins}}}$, obtained from a sequence of friction values $\mathcal{F}_{s}$ as
@@
Two hyperparameters must be defined, the number of bins ($n_{\text{bins}}$) and the limit values ($r_{\text{min}}$ and $r_{\text{max}}$). The latter can be determined from the historical data, by annotating the minimum and maximum values of friction observed. The number of bins to select is a trade-off between accuracy and complexity. A small number of bins may make distinguishing an anomaly impossible, while a large number will lead to a large histogram vector, which may complicate training and lead to overfitting. For our task, we select a value of $20$ bins, and later analyze the effect of this hyperparameter on the algorithm’s accuracy in Section V-2.
@@
The last decision to take is what range of friction values to include in the histogram. From the previous section, we recall that we obtain, for each subsystem, a collection of friction values paired with their corresponding configurations. Based on our knowledge of the anomaly, we can decide to include all friction values produced by the FSS in the histogram, or only those related to specific configurations (for example, only friction values obtained with $q=1$).
@@
The output of this final stage is the status of the RWA that generated $X$, given as $\hat{\theta}=(\hat{\theta}^{\mathrm{d}},\hat{\theta}^{\mathrm{v}},\{\hat{\theta}^{s}\}_{s=1}^{N_{\mathrm{s}}})$.
V Algorithm Tuning
Despite the complexity of the anomaly detection algorithm, most of its parameters are fairly straight–forward to tune, or have little impact on its results. There are, however, two parameters that require special care. The first is the bias weight ($W_{b}$), used in the changepoint detection stage to include the effect of the prior knowledge of the viscous friction coefficient. The second is the number of bins ($n_{\mathrm{bins}}$) that the SVM uses to classify anomalies related to FSS. In this section, we will study the effect of these two parameteres on the performance of the algorithm, and determine the best way to tune them.
V-1 Bias weight selection
Before deciding how to tune the bias weight, it is necessary to define what results we want to optimize. In the analysis of a changepoint detection algorithms, there are two main metrics of performance: the Missed Detection Rate (MDR), defined as the probability of not detecting a changepoint, and the Average Run Length (ARL), defined as the average number of non-anomalous points the algorithm goes through before detecting a false changepoint in the data [17]. Since it is always possible to improve one at the expense of the other (by decreasing or increasing the threshold $\text{GLR}_{\text{thr}}$), the challenge lies in achieving a high ARL with a low MDR.
To assess the performance of the changepoint detection, it is important to remember how the subsequent stages of the algorithm process changepoints. As mentioned in Section IV-C2, an undetected changepoint can invalidate the algorithm’s output while a false positive can be rejected during the assignment stage. This implies that while high ARL is beneficial, as it will lead to fewer false positives, a low MDR is required for the algorithm’s success.
For our analysis, we will fix the MDR to $0.1\%$ and determine the threshold value at which such sensitivity is achieved using the $\chi^{2}$ distribution. Then, we obtain the ARL with this threshold, which we call the $\text{ARL}^{0.1\%}$. A satisfactory value of ARL depends on the tolerance of the algorithm to false positives, and the number of points we process in a single window of measurements. In the TAS case, we analyze windows of around $80000$ measurements, so an ARL of around $10^{6}$ is a good target if we do not want to get more than $1$ false changepoint per window. Since the TAS dataset has no annotated changepoint locations, these results are obtained using our own simulations, in which we can place changepoints in a controlled way.
Back to the tuning of $W_{\mathrm{b}}$, we can run a parametric study of this coefficient to observe how it affects the $\text{ARL}^{0.1\%}$ for different window sizes, assuming that the changes of dry friction we are trying to detect have a magnitude of $3$ times the noise standard deviation. The results are shown in Figure 7, left plot. A higher window size $w$ leads to better algorithm results, as the effect of the noise is reduced when we use more points the w-GLR. However, increasing $w$ is not always feasible, because it is limited by the minimum distance we expect to see between changepoints. We can observe how the $W_{\mathrm{b}}$ improves the algorithm’s performance for different window sizes. As an example, using a window size of $20$, we can improve the $ARL^{0.01\%}$ from $10^{4}$ to $10^{10}$ by increasing $W_{\mathrm{b}}$ from $10^{-8}$ to $10^{-4}$. The bias term is, then, a good way of improving the performance of the changepoint detection when the window size is constrained, and it seems from this first analysis that $W_{\mathrm{b}}$ should be set to be arbitrarily large.
However, in the previous plot, we assumed that our estimated friction $\tilde{f}^{\mathrm{v}}$ is equal to the actual viscous friction $f^{\mathrm{v}}$. Unfortunately, this may not always be the case, since the latter is not known in advance. Considering the effect of the bias $\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}}$ is more complex: when the error is not zero, the w-GLR no longer follows a $\chi^{2}$ distribution. We can, however, estimate the effect of this error to be in the order of $W_{\mathrm{b}}(\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}})^{2}$. As the original distribution of the w-GLR is of order unit, we should ensure that $W_{\mathrm{b}}(\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}})^{2}\lll 1$ so the error does not have a noticeable effect. Using $W_{\mathrm{b}}=10^{-4}$, this implies that we need $|\tilde{f}^{\mathrm{v}}-f^{\mathrm{v}}|\lll 10^{2}$, an easily achievable condition since the viscous friction coefficient is of order $1$.
This gives a simple rule for tuning $W_{b}$: if the window size is large enough (more than 50 points), it can be set to 0. If not, a high value can be set to improve the results, but keeping in mind that if the prior of the viscous friction is not reliable, the performance may degrade and produce too many false positives. For our implementation of the algorithm, we will use $W_{b}=10^{-4}$.
V-2 Number of bins
The number of bins determines how the friction coefficients estimated for each FSS are translated into a vector that can be fed into a SVM to classify an anomaly. A small number of bins is desirable as it reduce the number of dimensions the SVM has to work with. However, it may also lead to loss of granularity in the transition from friction values to histograms, which can make the classification perform worse. As with the bias weight, we run a parametric test using different number of bins for classifying the anomaly in one of the FSS of the TAS dataset. The obtained misclassification rates for training and validation datasets is shown in Figure 8.
The plot shows an expected decrease in the train loss as the number of bins increases, since the dimension of the classification increases with it. However, many bins lead to overfitting, as training captures only the specific shape of the datapoints used in training, and the validation loss increases. This shows a good level of robustness of the algorithm to the number of bins, which can again be attributed to the use of SVMs for the classification. Since we see that the validation misclassification ration plateaus after around $40$ bins, we used this number in our algorithm.
VI Experimental Results